OpenStudy (anonymous):

Tony makes sandwiches for a lunch counter . This morning he made 50 beef and cheese sandwiches , 75 Ham and cheese sandwiches , And 16 Tuna sandwiches. The Boss wants 150 sandwiches how many more does he have to make for the Boss A(9 or B(19 or C(29 or D(39 will medal and fan

OpenStudy (anonymous):

@maddiegirl

OpenStudy (anonymous):

@dman2597

OpenStudy (anonymous):

@cutecaligirl_

OpenStudy (anonymous):

A is the answer because if you add 50+75+16=141+9=150 so your answer is A ,9